OpenStudy (anonymous):

Choose the correct simplification of (x2y4)3(3y2)2.

OpenStudy (anonymous):

a. 6x5y11 b. 9x5y11 c. 6x6y16 d. 9x6y16

OpenStudy (anonymous):

9x6y16

OpenStudy (anonymous):

so d

jhonyy9 (jhonyy9):

know you this 2a^2 *a^3 = ?

jhonyy9 (jhonyy9):

a^2 * a^3 = a^(2+3) = a^5 ok ?

jhonyy9 (jhonyy9):

(a^2)^3 = a^(2*3) = a^6 hope you understand it now sure easy

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (anonymous):

Thanks

jhonyy9 (jhonyy9):

yw